OpenStudy (anonymous):

function help given f(x) = 4√-x find: 1. f(4-x) 2. f(4-x^2) 3. f(4-x^4) 4. f(4x- x^2) 5. f(-x^4 - 4x^2) pls help me for my H.w

OpenStudy (anonymous):

plug in (4-x) in in place of x in function

OpenStudy (anonymous):

what is the answer for no.1?

OpenStudy (anonymous):

plug in (4-x) in in place of x in function and you will find out

OpenStudy (anonymous):

ic so in given f(4-x) = (4-x)? is this right ?

OpenStudy (anonymous):

F[x]=\[4 \sqrt{-x}\] plug in (4-x) in in place of x in function \[4 \sqrt{-(4-x)}\]

OpenStudy (anonymous):

ok i get it so after that what is next? example no.1

OpenStudy (anonymous):

Just Simplify it; this is no.1

OpenStudy (anonymous):

thank you so after wards for no.2 no.3 i just give square root right?

OpenStudy (anonymous):

discrete math function